# SpecExec: Massively Parallel Speculative Decoding for Interactive LLM Inference on Consumer Devices

### v2

@@ Abstract
As large language models gain widespread adoption, running them efficiently becomes crucial. Recent works on LLM inference use speculative decoding to achieve extreme speedups. However, most of these works implicitly design their algorithms for high-end datacenter hardware. In this work, we ask the opposite question: how fast can we run LLMs on consumer machines? Consumer GPUs can no longer fit the largest available models (50B+ parameters) and must offload them to RAM or SSD. When running with offloaded parameters, the inference engine can process batches of hundreds or thousands of tokens at the same time as just one token, making it a natural fit for speculative decoding. We propose SpecExec (Speculative Execution), a simple parallel decoding method that can generate up to 20 tokens per target model iteration for popular LLM families. It utilizes the high spikiness of the token probabilities distribution in modern LLMs and a high degree of alignment between model output probabilities. SpecExec takes the most probable tokens continuation from the draft model to build a “cache” tree for the target model, which then gets validated in a single pass. Using SpecExec, we demonstrate inference of 50B+ parameter LLMs on consumer GPUs with RAM offloading at 4–6 tokens per second with 4-bit quantization or 2–3 tokens per second with 16-bit weights.  [^1] The implementation code is available at https://github.com/yandex-research/specexec
@@ 1 Introduction
Open-access large language models (LLMs), such as Llama 3 [1] and Mistral [2], have become increasingly capable in the past years, and their adoption has grown dramatically. Although these models are openly available, users who are interested in running these models on consumer-grade GPUs (for example, due to privacy or cost reasons) face important challenges. Many open-access LLMs are too large to fit on consumer GPUs, which necessitates offloading them onto CPU RAM to perform inference. Given the limited memory bandwidth between CPU and GPU, as well as the fact that all model parameters must be transferred to the GPU for the LLM to generate each new token, offloading is extremely slow and bandwidth-bound. For example, generating a single new token using Llama 2-70B in 16 bit with offloading on an RTX 3090 GPU takes at least 4.5 seconds [^2] Assuming PCIe-4.0 and at least 140GB of DDR5 RAM with an efficient offloading implementation..
@@
A recent line of work that aims to speed up LLM inference is speculative decoding [3, 4], which uses a small draft model to predict the next tokens and a larger target model to verify which of those tokens to accept in parallel. Although speculative decoding is a promising direction, the speedups that existing methods can attain in the offloading setting are relatively modest. While studying existing approaches [3, 5, 6], we discovered that these methods do not scale well with the draft model token budget. In particular, as shown in Figure 1 (left), the number of tokens accepted by the target model is empirically upper bounded (by $\approx$10 for this model and dataset combination) regardless of the number of speculated tokens. In turn, methods that scale better with more draft tokens [7] rely on static tree structures that may not be optimal for every setting, as they require dedicated tree optimization for every change in the text domain, generation hyperparameters, and the hardware setup.
In this work, we aim to improve the effectiveness of speculative decoding for running large language models on consumer hardware with RAM offloading. We propose SpecExec, a speculative decoding method that addresses the performance, flexibility and scalability issues of prior methods. SpecExec [^3] We chose this name because our method directly applies speculative execution to LLM inference. The draft model “guesses” which token prefixes the target model will need to continue, and then the target model computes distributions of continuations with a single forward pass on the speculated prefix tree. adopts a powerful draft model to deterministically [^4] In contrast, speculative sampling requires stochastic generation of the draft tree using draft probabilities. construct a large draft tree that covers the most likely continuations of the prefix using a parallel search algorithm. We then apply a simple verification algorithm that views this tree as a cache of potential continuations and validates it with the target model in a single pass.
@@
    Using our observations, we designed a system that can run Llama 2-70B or comparable models interactively at 4-6 tok/s using 4-bit quantization or 2-3 tok/s with 16-bit weights on consumer GPUs with offloading, with 10-18x speedups vs sequential inference on the same hardware.
@@ 2.1 Speculative Decoding
In this study, we extend a family of algorithms for speculative decoding of autoregressive LLMs [8, 3, 4]. These algorithms generate tokens in two phases: drafting and verification.
During the drafting phase, the algorithm generates a candidate sequence of tokens by sampling from a small draft model $P(x_{t+1}|x_{0:t},\theta_{\text{draft}})$. In turn, the verification stage leverages the target model $P(x_{t+1}|x_{0:t},\theta_{\text{main}})$ to verify these draft sequences and accept all tokens that have passed the verification. The probability of accepting a token is chosen in a way that preserves the output distribution of sequential sampling from the original LLM [3, 4, 6]. A key advantage of speculative algorithms is that the main model can verify all draft tokens in parallel, which is more efficient than sequentially generating one token at a time.
Subsequent works in speculative decoding extend this idea in several directions, including generating multiple draft sequences or draft trees, using multiple draft models, and finetuning the draft models to improve generation speed [5, 10, 11]. Another line of follow-up studies explores alternative sources for the draft model: namely, [12] uses a subset of main model layers to produce a draft, [13] retrieves draft sequences from a search index, and [14] uses multiple stages of speculation. Leveraging these techniques, practitioners built efficient implementations for fast LLM inference [5, 15].
In our analysis, we focus on speculative decoding algorithms that support sampling from the target model and guarantee identical sample probabilities vs standard generation. The rationale for our choice is that most popular LLM applications (such as chat assistants) require stochastic sampling to introduce variability into their responses. This focus rules out several algorithms that only support greedy inference [9, 10]. Still, most works on speculative decoding fit within that criterion.
@@ 2.2 Parameter Offloading
Offloading has emerged as a crucial technique for leveraging GPUs with insufficient on-chip memory to hold entire models, to perform training or inference with these models. When using offloading, the model parameters are typically stored in CPU RAM, and layers are sequentially loaded onto the GPU when they are needed during a forward or backward pass  [16, 17, 18]. This method is most effective when the model is able to process many tokens in parallel (e.g., during training or large-batch inference), where the overhead of transferring layers between CPU and GPU is amortized over many computations  [19, 20]. However, it is less effective for small-batch inference, because generating each new token requires transferring the entire model from CPU to GPU, which is heavily bottlenecked by the memory bandwidth (even with PCIe 4.0).
@@ 2.3 Running LLMs on Consumer Devices
While our observations are not specific to any particular LLM, we focus on a practical case of running modern instruction-tuned models such as Llama-2-70B-chat [21] and Mixtral 8x7B [22]. To better estimate the target hardware setups, we study communities dedicated to running large models locally, such as [23]. A popular [^5] Based on desktop setups (from A, B, C, D, etc.), as well as popular hardware guides such as [24] hardware configuration for running those models locally is a desktop or a cloud instance with a single consumer-grade GPU [^6] For example, RTX 4060 or 4090 desktops, T4 or A2 VMs. with 12–24 GB VRAM, 4–8 CPU cores, 32–64 GB RAM, and a PCIe 3.0 or 4.0 x16 bus between CPU and GPU. Another popular setup is devices without a dedicated GPU, such as MacBooks with an ARM-based CPU, 16 GB RAM, and an SSD. While this survey may not be fully representative, it reveals popular setups that are not targeted by most speculative decoding research.
Running the largest language models in this setup requires either extreme compression or offloading. While it is possible to fit 70B+ LLMs into consumer GPUs by compressing them to 1.5–2 bits per parameter [25, 26], doing so causes such quantization errors that defeat the purpose of running large models [27, 26]. Thus, practitioners with consumer-grade hardware may find it optimal to run 50B+ models with mild (e.g., 4-bit) quantization and offload model parameters from GPU to RAM, or to SSD [18].
@@ 3 Preliminary analysis
Speculative decoding with offloading benefits from the fact that it is more efficient to process tokens in parallel than sequentially. In conventional inference, this is caused by the higher arithmetic intensity of GPU processing [^7] Parallel matrix multiplications do more useful computations per memory access for many tokens.. With offloading, there is a different bottleneck — loading model parameters from RAM (or SSD). Since modern offloading engines can dispatch model parameters in parallel with computation, the total processing time is the maximum of the time to load all parameters and the total computation time. In preliminary experiments (see Figure 1, right), we found that 70B models running on a consumer desktop can process hundreds or thousands of tokens within nearly the same time as just a single token.
This leads us to a question: how does speculative decoding perform when given hundreds to thousands of draft tokens? As shown in concurrent work [7], speculative decoding algorithms with single or multiple sequences, like SpecInfer, are effectively upper-bounded in the number of accepted tokens as the speculation budget grows. This is confirmed by our observations (see Figure 1, left), where even the more powerful Llama-2 7B saturates.
In regular GPU inference, using 7B draft models would be impractical, as the drafting steps would take too long. However, in our setting, large draft models can be justified because each offloaded forward pass takes significantly more than a second (see Figure 1, right). This runs contrary to a popular intuition in speculative decoding that favors smaller draft models [5, 10].
We also observe that sampling from modern LLMs often results in a few high-probability tokens that add up nearly to a probability of 1 (see Figure 2). If we can find these tokens using the draft model, we can construct a draft that will be accepted with a similarly high probability. In preliminary experiments for 70B models, we found that running beam search with a capable draft model (e.g., LLaMA-2 7B) can recover many of these high-probability tokens. Unfortunately, this kind of deterministic search is incompatible with speculative decoding for stochastic sampling, which called for alternative validation method.
Figure 2: 70B model cum. probability of top likely tokens vs draft model choice, OASST1 dataset.
@@ 4.1 Speculative Execution
As we observed in Section 3, high-probability continuations of large models are concentrated in a few tokens, and offloading benefits from running target model on hundreds or thousands of tokens. To use these observations, we formulate an alternative, simpler speculative decoding strategy. Unlike speculative decoding, SpecExec (short for "Speculative Execution") does not propose a new sampling procedure: it runs normal (sequential) sampling while trying to “guess” which probabilities will be needed during future steps and precomputing these samples in parallel. This is similar to speculative execution [31] in modern CPUs that predict which operations should be computed ahead of time to better utilize the compute cycles.
More specifically, whenever SpecExec uses target model probabilities, it looks them up in a speculative “cache”. If it encounters a token that is not in the cache, it queries the target model for that token and simultaneously computes probabilities for $B$ potential future tokens chosen with the draft model. If the draft model can guess the likely next tokens accurately enough, the algorithm will be able to run multiple sampling iterations using these cached probabilities without querying the target model until it “exhausts the cache” and begins anew. A formal description of SpecExec is given in Algorithm 1.
@@
18:    $\texttt{cache}[x_{\text{prefix}}\oplus\{x_{i}\}]=\texttt{next\_probs}[x_{i}]$$\triangleright$ vector of probs for possible next tokens
@@
However, this advantage comes at the cost of lower acceptance rates for any individual token. Algorithm 1 accepts a token $x_{t}$ with probability $P(x_{t+1}|x_{0:t},\theta_{\text{target}})$, because accepting a token with SpecExec is equivalent to sampling that token from the target model distribution. Meanwhile, the original speculative decoding (for example, 5) accepts tokens with a higher probability ${P(x_{t+1}|x_{0:t},\theta_{\text{target}})}/{P(x_{t+1}|x_{0:t},\theta_{\text{draft}})}$.
For a small number of draft tokens (for instance, just one token), SpecExec is less effective than traditional speculative decoding. However, as we increase the number of draft tokens, speculative execution generates better-structured trees, which in practice leads to accepting more tokens for the same draft size: we verify this in Section 5.2.
Correctness. Next, we need to verify that SpecExec is equivalent to sequential sampling from the target model. Notably, unlike [3], SpecExec does not change the probabilistic sampling procedure. The difference between SpecExec and sequential decoding is that SpecExec computes some probabilities ahead of time, thus improving the GPU utilization in the case of offloading.
From a formal perspective, we rely on the fact that a speculative generation algorithm is equivalent to sequential sampling if it is locally equivalent in every node [5]; in other words, it samples from the same probabilities for every prefix in the draft tree. Since SpecExec explicitly samples from the same probabilities as the main model, this is true by definition.
The fact that SpecExec follows the same sampling procedure has another side effect. If we view SpecExec as a deterministic function that depends on a pseudo-random number generator as input, we can prove a stronger degree of equivalence. Namely, for every input random seed, SpecExec produces exactly the same outputs as sequential sampling with the same seed. In contrast, speculative decoding does not have this properly: it only guarantees correct overall sampling probabilities.
@@ 4.2 Search for the Optimal Draft Tree
As we discussed above, our algorithm uses the draft model to build a tree $\tau$ of likely future tokens for speculative caching. In this section, we describe how to find these tokens efficiently. From an optimization perspective, we seek to construct a tree that will lead to the highest expected number of generated (accepted) tokens. This problem can be solved by viewing the tree construction as the search for the set of nodes (i.e., tokens) that have the highest cumulative probability with respect to the target model. As we show in Appendix A, this search can be reduced to the single-source shortest path (SSSP) search problem that can be solved efficiently using a modified version of the Dijkstra’s algorithm [32], described formally in Algorithm 2.
@@
While Algorithm 2 is a rather special case of SSSP over a combinatorially large tree (the tree of all token sequences up to length $K$), the general SSSP problem is well studied in the computer science community (see Appendix B). Therefore, practitioners will be able to leverage existing algorithms to implement Speculative Execution for a broad range of setups, including GPUs, mobile CPUs, or distributed systems.
@@
7:  $H:=\textsc{PriorityQueue}(\{x:0\})$$\triangleright$ $x$ has priority 0; H orders by negative cum. log-prob.
@@
20:    $\texttt{topk}:=\textsc{selectBest}(\texttt{batch},\texttt{probs},\tau,K)$triangleright select best tokens by cumulative prob.

### v3

@@ Abstract
As large language models gain widespread adoption, running them efficiently becomes a crucial task. Recent works on LLM inference use speculative decoding to achieve extreme speedups. However, most of these works implicitly design their algorithms for high-end datacenter hardware. In this work, we ask the opposite question: how fast can we run LLMs on consumer machines? Consumer GPUs can no longer fit the largest available models and must offload them to RAM or SSD. With parameter offloading, hundreds or thousands of tokens can be processed in batches within the same time as just one token, making it a natural fit for speculative decoding. We propose SpecExec (Speculative Execution), a simple parallel decoding method that can generate up to 20 tokens per target model iteration for popular LLM families. SpecExec takes the most probable continuations from the draft model to build a “cache” tree for the target model, which then gets validated in a single pass. Using SpecExec, we demonstrate inference of 50B+ parameter LLMs on consumer GPUs with RAM offloading at 4–6 tokens per second with 4-bit quantization or 2–3 tokens per second with 16-bit weights.  [^1] The code is available at github.com/yandex-research/specexec.
@@ 1 Introduction
Open-access large language models (LLMs), such as Llama (Touvron et al., 2023) and Mistral (Jiang et al., 2023), have become increasingly capable in the past years, and their adoption has grown dramatically. Although these models are openly available, users who are interested in running these models on consumer-grade GPUs (for example, due to privacy or cost reasons) face significant challenges. Many open-access LLMs are too large to fit on consumer GPUs, which necessitates offloading them onto CPU RAM to perform inference. Given the limited memory bandwidth between the CPU and the GPU, as well as the fact that all model parameters must be transferred to the GPU for the LLM to generate each new token, offloading is extremely slow and bandwidth-bound. For example, generating a single token using Llama 2-70B in 16 bit with offloading on an RTX 3090 GPU takes at least 4.5 seconds [^2] Assuming PCIe-4.0 and at least 140GB of DDR5 RAM with an efficient offloading implementation..
@@
A recent line of work that aims to speed up LLM inference is speculative decoding (Leviathan et al., 2023; Chen et al., 2023a), which uses a small draft model to predict the next tokens and a larger target model to verify which of those tokens to accept in parallel. Although speculative decoding is a promising direction, the speedups that existing methods can attain in the offloading setting are relatively modest. While studying existing approaches (Leviathan et al., 2023; Miao et al., 2023; Sun et al., 2023), we discovered that these methods do not scale well with the draft model token budget. In particular, as shown in Figure 1 (left), the number of tokens accepted by the target model is empirically upper-bounded (approximately by 10 for this model and dataset combination) regardless of the number of speculated tokens. In turn, methods that scale better with more draft tokens Chen et al. (2024a) rely on static tree structures that may not be optimal for every setting, as they require tree structure optimization for every change in the text domain, generation parameters, and the hardware setup.
In this work, we aim to improve the effectiveness of speculative decoding for running large language models on consumer hardware with RAM offloading. We propose SpecExec, a speculative decoding method that addresses the performance, flexibility and scalability issues of prior methods. SpecExec [^3] We chose this name because our method directly applies speculative execution to LLM inference. The draft model “guesses” which token prefixes the target model will need to continue, and then the target model computes distributions of continuations with a single forward pass on the speculated prefix tree. adopts a powerful draft model to deterministically [^4] In contrast, speculative sampling requires stochastic generation of the draft tree using draft probabilities. construct a large draft tree that covers the most likely continuations of the prefix with a parallel search algorithm. We then apply a simple verification algorithm that views this tree as a cache of potential continuations and validates it with the target model in a single pass.
@@
    Using our observations and SpecExec, we design a system that can run Llama 2-70B or comparable models interactively at 4–6 tokens/second using 4-bit quantization or 2–3 tokens/second with 16-bit weights on consumer GPUs with offloading, with 10–18x speedups compared to sequential inference on the same hardware.
@@ 2.1 Speculative Decoding
In this study, we extend a family of algorithms for speculative decoding of autoregressive LLMs Stern et al. (2018); Leviathan et al. (2023); Chen et al. (2023a). These algorithms generate tokens in two phases: drafting and verification.
During the drafting phase, the algorithm generates a candidate sequence of tokens by sampling from a small draft model $P(x_{t+1}|x_{0:t},\theta_{\text{draft}})$. In turn, the verification stage leverages the target model $P(x_{t+1}|x_{0:t},\theta_{\text{main}})$ to verify these draft sequences and accept all tokens that have passed the verification. The probability of accepting a token is chosen in a way that preserves the output distribution of sequential sampling from the original LLM Leviathan et al. (2023); Chen et al. (2023a); Sun et al. (2023). A key advantage of speculative algorithms is that the main model can verify all draft tokens in parallel, which is more efficient than sequentially generating one token at a time.
Subsequent works in speculative decoding extend this idea in several directions, including generating multiple draft sequences or draft trees, using multiple draft models, and finetuning the draft models to improve generation speed Miao et al. (2023); Liu et al. (2023); Xu et al. (2023). Another line of follow-up studies explores alternative sources for the draft model: namely, self-speculative decoding uses a subset of main model layers to produce a draft Zhang et al. (2023), REST retrieves draft sequences from a search index He et al. (2023), and staged speculative decoding uses multiple levels of speculation Spector and Re (2023). Leveraging these techniques, practitioners have built efficient implementations for fast LLM inference (Miao et al., 2023; Cai et al., 2023). We refer the readers to survey papers for a more detailed coverage of speculative decoding methods Zhang et al. (2024a); Xia et al. (2024a).
In our analysis, we focus on speculative decoding algorithms that support sampling from the target model and guarantee identical sample probabilities vs standard generation. The rationale for our choice is that most popular LLM applications (such as chat assistants) require stochastic sampling to introduce variability into their responses. This focus rules out several algorithms that only support greedy inference Fu et al. (2023); Liu et al. (2023). Still, most works on speculative decoding fit within that criterion.
@@ 2.2 Parameter Offloading
Another recent line of work explores running and training large models with limited accelerator memory by “offloading” their parameters to more abundant storage, such as RAM or even SSD (Pudipeddi et al., 2020; Ren et al., 2021; Alizadeh et al., 2023). This technique works by loading model parameters on the GPU when they are needed for computation. Since most deep learning models use layers in a fixed order, offloading can pre-dispatch the next layer’s parameters in the background.
This technique works particularly well when processing large batches of data, during training Pudipeddi et al. (2020); Ren et al. (2021) or large-batch non-interactive inference Aminabadi et al. (2022); Sheng et al. (2023), where each layer process multiple tokens each time it is loaded. In turn, when doing interactive inference, offloading works significantly slower than on-device inference. This is because interactive inference has to process one or few tokens at a time, and therefore spends most of the time waiting for the parameters to load.
@@ 2.3 Running LLMs on Consumer Devices
While our observations are not specific to any particular LLM, we focus on a practical case of running modern instruction-tuned models such as Llama-2-70B-chat Touvron et al. (2023) and Mixtral 8x7B Jiang et al. (2024). To better estimate the target hardware setups, we study communities dedicated to running large models locally, such as LocalLlama (2023). A popular [^5] Based on popular hardware guides such as Dettmers (2023) as well as setups examples (see A, B, C, D) hardware configuration for running those models locally is a desktop or a cloud instance with a single consumer-grade GPU [^6] For example, RTX 4060 or 4090 desktops, T4 or A2 VMs. with 12–24 GB VRAM, 4–8 CPU cores, 32–64 GB RAM, and a PCIe 3.0 or 4.0 x16 bus between CPU and GPU. Another popular setup is devices without a dedicated GPU, such as MacBooks with an ARM-based CPU, 16 GB RAM, and an SSD. While this survey may not be fully representative, it reveals popular setups that are not targeted by most speculative decoding research.
Running the largest language models in this setup requires either extreme compression or offloading. While it is possible to fit 70B+ models into consumer GPUs by compressing them to 1.5–2 bits per parameter Chee et al. (2023); Tseng et al. (2023), doing so causes significant accuracy losses that defeat the purpose of running large models Dettmers and Zettlemoyer (2022); Tseng et al. (2023). Thus, practitioners with consumer-grade hardware may find it optimal to run 50B+ models with mild (e.g. 4-bit) quantization and offload parameters from GPU to RAM or SSD Alizadeh et al. (2023).
@@ 3 Preliminary analysis
Speculative decoding with offloading benefits from the fact that it is more efficient to process tokens in parallel than sequentially. In conventional inference, this is caused by the higher arithmetic intensity of GPU processing [^7] Parallel matrix multiplications do more useful computations per memory access for multiple tokens.. With offloading, there is a different bottleneck — loading model parameters from storage. Since offloading engines can dispatch model parameters in parallel with computation, the total processing time is the maximum of the time to load all parameters and the total computation time. In preliminary experiments (see Figure 1, right), we found that 70B models running on a consumer desktop can process thousands of tokens within nearly the same time as just a single token.
This leads us to a question: how does speculative decoding perform when given hundreds to thousands of draft tokens? As shown in concurrent work Chen et al. (2024a), speculative decoding algorithms with single or multiple sequences, like SpecInfer, are effectively upper-bounded in the number of accepted tokens as the speculation budget grows. This is confirmed by our observations (see Figure 1, left), where the number of accepted tokens saturates even for the more powerful Llama-2 7B.
In regular GPU inference, using 7B draft models would be impractical, as the drafting steps would take too long. However, in our setting, large draft models can be justified because each offloaded forward pass takes significantly more than a second (see Figure 1, right). This runs contrary to a popular intuition in speculative decoding that favors smaller draft models Miao et al. (2023); Liu et al. (2023).
We also observe that sampling from modern LLMs often results in a few high-probability tokens that add up nearly to a probability of 1 (see Figure 2 for an illustration). If we can find these tokens using the draft model, we can construct a draft that will be accepted with a similarly high probability. In preliminary experiments for 70B models, we found that running beam search with a capable draft model (e.g., Llama-2 7B) can recover many of these high-probability tokens. Unfortunately, this kind of deterministic search is incompatible with speculative decoding for stochastic sampling, which called for alternative validation method.
Figure 2: Llama-2 70B Chat model cumulative probability of most likely tokens compared to the draft model choice (all Llama draft models are chat versions), OASST1 dataset.
@@ 4.1 Speculative Execution
As we observed in Section 3, high-probability continuations of large models are concentrated in a few tokens, and offloading benefits from running target model on hundreds or thousands of tokens. To use these observations, we formulate an alternative, simpler speculative decoding strategy. Unlike speculative decoding, SpecExec (short for “Speculative Execution”) does not propose a new sampling procedure: it runs standard (sequential) sampling while trying to “guess” which probabilities will be needed during future steps and precomputing these continuations in parallel. This is similar to speculative execution (Lampson, 2006) in modern CPUs that predicts which operations should be computed ahead of time to better utilize the compute cycles.
More specifically, whenever SpecExec uses target model probabilities, it looks them up in a speculative “cache”. If it encounters a token that is not in the cache, it queries the target model for that token and simultaneously computes probabilities for $B$ potential future tokens chosen with the draft model, where $B$ is the batch size. If the draft model can guess the likely next tokens accurately enough, the algorithm will be able to run multiple sampling iterations using these cached probabilities without querying the target model until it “exhausts the cache” and begins anew. A formal description of SpecExec is given in Algorithm 1.
@@
18:    $\texttt{cache}[x_{\text{prefix}}\oplus\{x_{i}\}]=\texttt{next\_probs}[x_{i}]$$\triangleright$ probabilities of possible next tokens
@@
However, this advantage comes at the cost of lower acceptance rates for any individual token. Algorithm 1 accepts a token $x_{t}$ with probability $P(x_{t+1}|x_{0:t},\theta_{\text{target}})$, because accepting a token with SpecExec is equivalent to sampling that token from the target model distribution. Meanwhile, the original speculative decoding (for example, Miao et al. (2023)) accepts tokens with a higher probability ${P(x_{t+1}|x_{0:t},\theta_{\text{target}})}/{P(x_{t+1}|x_{0:t},\theta_{\text{draft}})}$.
For a small number of draft tokens (for instance, just one token), SpecExec is less effective than traditional speculative decoding. However, as we increase the number of draft tokens, speculative execution generates better-structured trees, which in practice leads to accepting more tokens for the same draft size; we verify this in Section 5.2.
Correctness. Next, we need to verify that SpecExec is equivalent to sequential sampling from the target model. Notably, unlike Leviathan et al. (2023), SpecExec does not change the probabilistic sampling procedure. The difference between SpecExec and sequential decoding is that SpecExec precomputes some probabilities, thus improving the GPU utilization in the case of offloading.
From a formal perspective, we rely on the fact that a speculative generation algorithm is equivalent to sequential sampling if it is locally equivalent in every node Miao et al. (2023); in other words, it samples from the same probabilities for every prefix in the draft tree. Since SpecExec explicitly samples from the same probabilities as the main model, this is true by construction.
The fact that SpecExec follows the same sampling procedure has another side effect. If we view SpecExec as a deterministic function that depends on a pseudo-random number generator as input, we can prove a stronger degree of equivalence. Namely, for every seed value of the random number generator, SpecExec produces exactly the same outputs as sequential sampling with the same seed. In contrast, speculative decoding does not have this properly, as it only guarantees correct probabilities for the overall generation procedure.
@@ 4.2 Search for the Optimal Draft Tree
As we discussed above, our algorithm uses the draft model to build a tree $\tau$ of likely future tokens for speculative caching. In this section, we describe how to find these tokens efficiently. From an optimization perspective, we seek to construct a tree that will lead to the highest expected number of generated (accepted) tokens. This problem can be solved by viewing the tree construction as the search for the set of nodes (i.e., tokens) that have the highest cumulative probability with respect to the target model. As we show in Appendix A, this search can be reduced to the single-source shortest path (SSSP) search problem that can be solved efficiently using a modified version of the Dijkstra’s algorithm Dijkstra (1959), described formally in Algorithm 2.
@@
For a visual high-level representation of the SpecExec algorithm, see Appendix B.
While Algorithm 2 is a rather special case of SSSP over a combinatorially large tree (the tree of all token sequences up to length $K$), the general SSSP problem is well studied in the computer science community (see Appendix C). Therefore, practitioners will be able to leverage existing algorithms to implement Speculative Execution for a broad range of setups, including GPUs, mobile CPUs, or distributed systems.
@@
7:  $H:=\textsc{PriorityQueue}(\{x:0\})$triangleright $x$ has priority 0; H is ordered by negative cumulative log-probabilities
@@
20:    $\texttt{topk}:=\textsc{selectBest}(\texttt{batch},\texttt{probs},\tau,K)$triangleright select best tokens by cumulative probability